OpenStudy (anonymous):

Whats the anti derivative of (1/(x^4)) ?? Need Help

OpenStudy (amistre64):

x^-4 x^-3 ---- i think -3

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

\(\frac{-1}{3x^3}+C\) lol

OpenStudy (amistre64):

it looks to work like a derivative rule....

OpenStudy (anonymous):

\[\int\limits_{}^{}x^{n}dx=x^{n+1}/(n+1)+C\] here n=-4

OpenStudy (amistre64):

1/x^4 = -1/3x^3 ....close..

OpenStudy (anonymous):

so you dont brink the x ^4 to the numerater?

OpenStudy (anonymous):

it's in the denominator so to bring it to the numerator you switch the sign of the exponent

OpenStudy (anonymous):

ya i know that , but how is the one negative?

OpenStudy (amistre64):

\[\frac{x^{-4+1}}{-4+1}\]

OpenStudy (amistre64):

\[\frac{x^-{3}}{-3}\]

OpenStudy (amistre64):

\[-\frac{1}{3x^3}\]

OpenStudy (anonymous):

thank you